OpenStudy (anonymous):

need to solve for x and can't do it! help! -1/x^2+1/x=0

OpenStudy (amistre64):

-1+x ------ = 0 x^2 x = 1

OpenStudy (anonymous):

\[-1/x^2+1/x=0\]\[-1/x^2+x/x^2=0\]

OpenStudy (dumbcow):

shortcut way, add 1/x2 to other side flip both the fractions x=x^2 x=1

OpenStudy (anonymous):

\[-1+x=0, x=1\]